Below are the comparative balance sheets for 2016 and 2017 as well as the income statement for 2017 for Dynamite Sales company.

2017

2016

Income statement

Sales revenue

$845,000

Cost of goods sold

520,000

Gross profit

325,000

Operating expenses

110,000

Pretax income

215,000

Income tax

47,300

Net income

$167,700

Balance sheet

Cash

$7,200

$5,620

Accounts receivable (net)

39,400

36,000

Merchandise inventory

14,000

18,000

Prepaid-expenses

500

400

Fixed assets (net)

210,000

180,000

$271,100

$240,020

Accounts payable

$19,000

$18,000

Income taxes payable

2,000

1,000

Bonds payable

94,000

90,000

Common stock

100,000

100,000

Retained earnings

56,100

31,020

$271,100

$240,020

*Assume the following. Interest expense is $8,000. The market price of stock on December 31, 2015 is $26.00. The income tax rate is 22%. The average number of shares outstanding over 2015 is 10,000 shares.

Using this information, calculate and show work for each of the following ratios (show as decimals to four places – do not show as percentages):

Profit margin

Return on assets

Return on equity

Earnings per share

Price/earnings ratio

Current ratio

Quick ratio

Acid test ratio

Debt-to-equity ratio

Accounts receivable turnover

A/R days until received
Answer #1
  1. Profit margin= Net Income/Net Sales=$167700/$845000

                                 =0.1985

  1. Return on assets=65.62%

Assets in the beginning (2016)=$ 240020

Assets at the end(2017)=$271100

Average assets= (opening assets+closing assets)/2=$511120/2=$ 255560

Return on assets=Net income/average assets=$167700/$255560

                                       =0.6562

  1. Return on equity=1.68

Return on equity=net income/shareholder’s equity

                              =$167700/$100000=1.677

  1. Earnings per share=$16.77 per share

Earnings per share=Net income/outstanding shares at the end

                                      =$167700/10,000 shares=$16.77

  1. Price/Earnings ratio=1.55

Price/Earnings ratio=Market Price/Earnings per share

                                  =$26/$16.77=1.5504